The ancient Greeks thought that the
most pleasing shape for a rectangle was one for which the
ratio of the length to the width was approximately 8 to 5,
the golden ratio. If the length of a rectangular painting is
2 ft longer than its width, then for what dimensions would
the length and width have the golden ratio?

Answer by checkley77(12844)

(x+2)/x=8/5
8x=5(x+2)
8x=5x+10
8x-5x=10
3x=10
x=10/3
x=3.3333 ans. for the width.
3.3333+2=5.3333 ans. for the length.
Proof:
5.3333/3.3333=8/5
1.6~1.6